Q2. Solve the following inequality: 2r -1 3 A rectangular field has a perimeter of 104 meters. The length of the field is 12 meters more than its width. Find the length and the width of this field. Q3.
Answer #1

2) \frac{2x - 1}{-3} < -5

or, 2x - 1 < 15

or, 2x < 16

or, x < 8

3) Let the width = x meter

Length = (x + 12) meter

The perimeter of the rectangular field is 104

2(x + x + 12) = 104

or, 2(2x + 12) = 104

or, 2x + 12 = 52

or, 2x = 40

or, x = 20

So the width is 20 meter

And the length = 20 + 12 = 32 meter